Your client is 30 years old and wishes to retire at age 60. At the time, your client quashes to have saved $2,000,000. You advise the client to set aside money every year for the next 30 years. This money will be invested in a fund that you believe will average 12% each year for the next 30 years.

A) How much will your client need to set aside for each payment into the fund in order to accumulate the $2,000,000 total?

B) If your client’s life expectancy is age 80, how much can your client withdraw each year for the 20 year retirement? The cost of money is 12%.
Answer #1

A:  Future Value of annuity = A* ((1+rate)^n-1)/rate

2000000= A*((1+12%)^30-1)/12%

2000000=A*241.3327

A= 82873.15

Amount to set aside every year = $8287.32

B: PV of annuity = Annuity*(1-1/(1+rate)^number of terms)/rate

2000000=A*(1-1/1.12^20)/0.12

2000000=A*7.4694

A= 267757.56